OpenStudy (anonymous):

Given f(x) = 2x^2 + 2x and g(x) = x + 1, find the function (f/g)(x).

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

make a fraction bar, put on over the other

OpenStudy (misty1212):

\[\frac{f}{g}(x)=\frac{f(x)}{g(x)}\]

OpenStudy (anonymous):

Okay I just did!

OpenStudy (misty1212):

you got that? just put \(2x^2+2x\) on top of \(x+1\) they may want you to factor and cancel, although that is wrong we can do it anyway if you like

OpenStudy (misty1212):

\[\frac{2x^2+2x}{x+1}=\frac{2x(x+1)}{x+1}=2x\] if \(x\neq -1\)

OpenStudy (anonymous):

Thanks that makes so much more sense

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

c: